OpenStudy (toxicsugar22):

calcute the value of [-OH] from the given [H3O+] in each solution and label the solution as acidic or basic A) [H3O+]=10^-3 M; B) [H30+]=10^-9 M.

rvc (rvc):

@Somy

rvc (rvc):

i gtg

OpenStudy (somy):

this is full question?

OpenStudy (toxicsugar22):

yes

OpenStudy (somy):

@Kainui sweetheart :* help this kid plz i need to leave :3

OpenStudy (toxicsugar22):

see if you can help me

OpenStudy (anonymous):

it so easy

OpenStudy (triciaal):

is that all of the question? what solution? Chemistry was fun but I really don't remember much, sorry

OpenStudy (anonymous):

we know that [H3O+] * [OH-] = \[10^{-14}\] and also PH = 14 - POH or PH= -Log (H3O+) and POH = -Log (OH-) us i know .......... welcome

OpenStudy (toxicsugar22):

OK so i dont see how you get the solution

OpenStudy (toxicsugar22):

i dont get it

OpenStudy (triciaal):

@joannablackwelder will you please help

OpenStudy (triciaal):

@preetha will you be able to help?

OpenStudy (anonymous):

for the solution A is an acide PH= 3 <7 and B is basic PH= 9 > 7

OpenStudy (triciaal):

@Tech1 will you please help? your knowledge is more recent? I have forgotten a lot of this.

OpenStudy (toxicsugar22):

for the place where we put the answer is A) _____*10___M WHAT DO I FILL THESE WITH BY

OpenStudy (anonymous):

For the solution A we have : [H3O+]=10^-3 M; ====> Ph = 3 =====> POH = 14-3 = 11 =====> \[[OH-] = 10^{-11}\] ======> there are a little of OH- in A so is an Acid solution

OpenStudy (anonymous):

For Solution B we have : [H30+]=10^-9 M ===> PH=9 ===> POH = 14-9 = 5 ====>\[[OH−]=10^{-5}\] ====so====> in B there are some of OH- so its a basic solution

OpenStudy (toxicsugar22):

is it 1.00*10^-11

OpenStudy (toxicsugar22):

and B

OpenStudy (anonymous):

same

OpenStudy (toxicsugar22):

1.0

OpenStudy (anonymous):

\[1*10^{-5}\]

OpenStudy (toxicsugar22):

THANKS

OpenStudy (anonymous):

welcome .....(who?)

OpenStudy (toxicsugar22):

CAN YOU HELP WITH ONE MORE